# ♪  Something Just Like TRuST ♪ thanks: An allusion to Coldplay’s “Something Just Like This.” : Toxicity Recognition of Span and Target

### v1

@@ 3.2 Toxicity and Target Group Labels
Table 1 shows the total count for each higher- or lower-level target group in the human-annotated subset, along with the proportion of these that are labeled toxic. Most target groups are ethnicity-based; the least frequent target group is for "disability." In general, we have almost a balanced dataset for toxic versus non-toxic content (47.37% of the examples are toxic), but the rate varies greatly within each group; for example, 75% of the examples for the social group "black" are toxic, while only 33% are for "native american."
3.3 Span Labels
@@
The total number of examples with annotation of the toxic span is 5,506 (47.89% of 11,498, per Table 1), where some examples have multiple spans. To calculate descriptive statistics on span tokens, we first apply stemming. We merge the strings “ni**a” and “ni**er”. The mean length of spans is 1.91 words, the median is 1, and the maximum is 11. In 33% of cases with span annotation, the span constituted the entire sentence. There are 7065 spans in total, among which 1334 are unique. The histogram in Figure 1 shows the span tokens occurring more than once is highly skewed. We show the 25 most common span tokens, with “ni**er”, “f**k”, “stupid”, “idiot”, and “b**ch” at the top. Some words are specific to particular groups such as “black” or “kike” but others such as “kill” or “stupid” are used more generally.
@@
Model Accuracy Precision Recall F1 PerspectiveAPI 0.50 0.48 0.77 0.59 RoBERTa 0.79 0.76 0.82 0.79 BERT 0.79 0.77 0.79 0.78 GPT4o 0.75 0.70 0.85 0.77 Sonnet 0.77 0.72 0.83 0.77 Llama70b 0.77 0.71 0.86 0.78 Llama8b 0.73 0.66 0.88 0.76 Reasoning and CoT D. Llama70b 0.75 0.69 0.84 0.76 D. Llama8b 0.70 0.63 0.90 0.74 o4-mini 0.78 0.71 0.90 0.79 GPT4o-cot 0.74 0.73 0.71 0.72 Llama8b-cot 0.73 0.67 0.84 0.75 Llama70b-cot 0.75 0.72 0.75 0.74 Sonnet-cot 0.73 0.73 0.67 0.70
Table 2: Toxicity detection results. D. models are R1 distilled versions.
4 Experiments
@@
Our experiments utilize the human-annotated data of 11,498 examples divided into validation, training and test as follows. The validation set is 495 examples, with randum selection plus enforcing a minimum of 5 examples per target group. The randomly selected test set has 982 examples. We compare performance of multiple baselines on the three tasks of detection of toxicity, target social group, and span detection for toxic examples. We first compare PLMs versus zero-shot LLMs on each task in turn, including a custom model in the case of toxicity detection, followed by a subsection on whether prior knowledge of target social group improves toxicity detection. Then we compare these results to the use of reasoning models and in-context learning. We report accuracy, precision, recall and F1 for the first two tasks, but omit accuracy for toxic span detection, as the exact matching needed for accuracy is less informative than the breakdown of F1 into recall and precision on toxic span words. The final subsection presents an error-analysis.
@@
PLM baselines for toxicity and target social group use BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019) with linear classifier layers, and for span prediction we use SpanBERT Joshi et al. (2020) (see appendix B). We include four LLMs: GPT4o OpenAI et al. (2024), Claude 3.7 Sonnet Anthropic (2025), and Llama3.1 (70b and 8b) Grattafiori et al. (2024) (see Appendix E for the prompt). We use temperature=0, set the seed, and use the default values for others for determinism, although it is not guaranteed Atil et al. (2025). For toxicity, we also include PerspectiveAPI Lees et al. (2022), a neural network that provides a probability that an example is toxic. We used the validation set to identify the best probability for the binary class cutoff, which was 0.20.
@@ 4.1 Toxicity
The toxicity results in the top of Table 2 show that fine-tuned PLMs perform slightly better than the LLMs. Except for Llama8b, LLMs perform similarly. Surprisingly, PerspectiveAPI’s accuracy is random. All models have higher recall which is preferable here, where false negatives are worse than false positives.
Model Accuracy Precision Recall F1 BERT 0.76 0.68 0.82 0.72 RoBERTa 0.73 0.63 0.81 0.70 GPT4o 0.75 0.67 0.78 0.70 Sonnet 0.74 0.62 0.75 0.67 Llama70b 0.65 0.55 0.68 0.58 Llama8b 0.48 0.15 0.15 0.15 Reasoning and CoT o4-mini 0.72 0.58 0.63 0.59 D. Llama70b 0.69 0.24 0.27 0.25 D. Llama8b 0.61 0.17 0.17 0.17 GPT4o-cot 0.75 0.68 0.75 0.70 Llama8b-cot 0.40 0.07 0.06 0.06 Llama70b-cot 0.68 0.50 0.58 0.53 Sonnet-cot 0.75 0.62 0.70 0.65
@@
Model Accuracy Precision Recall F1 BERT 0.80 0.77 0.82 0.79 RoBERTa 0.78 0.74 0.81 0.77 GPT4o 0.66 0.64 0.65 0.64 Sonnet 0.71 0.65 0.70 0.66 Llama70b 0.57 0.46 0.48 0.44 Llama8b 0.49 0.07 0.07 0.07 Reasoning and CoT o4-mini 0.67 0.45 0.47 0.45 D. Llama70b 0.64 0.17 0.18 0.17 D. Llama8b 0.56 0.07 0.07 0.07 GPT4o-cot 0.68 0.56 0.61 0.58 Llama8b-cot 0.44 0.05 0.05 0.05 Llama70b-cot 0.57 0.31 0.31 0.31 Sonnet-cot 0.70 0.55 0.60 0.57
@@ 4.2 Target Social Group
Similar to Zampieri et al. (2023), for two baselines we train a neural network with a linear classifier layer on top of BERT or RoBERTa encoder. As above, we also test the same SOTA LLMs (cf. appendices F and G for the prompts).
Similar to toxicity detection results, fine-tuning a model slightly outperforms LLMs, as seen for the fine-grained target results in Table 3. However, Llama models do not perform as well as GPT4o and Sonnet. The F1 for Llama8b is especially low, showing an inability to understand social targets of toxicity. Table 4 shows the results for the higher-level target groups. Interestingly, LLMs perform worse at this task than the fine-grained target groups (9% decrease for GPT4o, 3% decrease for Sonnet etc.). Confusion matrices for GPT4o and Sonnet (the two top-performing LLMs), show that GPT4o mixes “other” and “ethnicity” predictions with “no target”, and Sonnet mixes “no target” with “ethnicity”. Additionally, both mix “ethnicity” with “other”. This indicates “other” as a high level target is more confusing than the fine-grained targets.
Model Precision Recall F1 SpanBERT 0.72 0.71 0.70 GPT4o 0.55 0.79 0.65 Sonnet 0.66 0.45 0.53 Llama70b 0.66 0.22 0.33 Lama8b 0.48 0.48 0.48 Reasoning and Cot o4-mini 0.63 0.40 0.49 D. Llama70b 0.45 0.87 0.59 D. Llama8b 0.43 0.58 0.4 Sonnet-cot 0.63 0.5 0.56 GPT4o-cot 0.55 0.63 0.59 Llama70b-cot 0.68 0.46 0.54 Llama8b-cot 0.68 0.16 0.26
@@
(a) Toxicity few-shot results.
(b) Target group few-shot results.
(c) Higher Target Group Few-Shot Results.
(d) Toxic Span Few-Shot Results.
Figure 2: Few-Shot Comparison Figures
@@ 4.3 Toxic Span
Table 5 shows that SpanBERT outperforms the LLMs. GPT4o’s F1 approaches SpanBERT’s, but the other LLMs do much worse.
4.4 Target Group and Toxicity Detection
We conduct experiments to assess whether prior knowledge about target social group could improve toxicity detection. For the RoBERTa/BERT-based models, we either add this information at the text level as “The target social group is <social group>” or at the embedding level. For LLMs, we use two types of prompting. In the first, we assign the social group persona to the model; combining persona with self-correction has been effective Xu et al. (2024). In the second, we include the social group in the text of the prompt. We find no differences between the methods (cf. Table 7 of the Appendix).
@@ 4.5 Reasoning Models
We also experiment with chain-of-thought (CoT) Wei et al. (2023) and the reasoning models o4-mini and R1 Distilled Llama70b/8b . Although reasoning helps in science or logic Jaech et al. (2024); DeepSeek-AI (2025); Zhang et al. (2024); Wei et al. (2023), it does not improve detection of toxicity or target group, and gives mixed results on span prediction. CoT helps Llama70b and Sonnet, but not Llama8b and Gpt4o. Reasoning usually increases recall, which indicates that models predict more tokens less precisely.
Figure 3: Toxicity detection accuracy by target group
@@ 4.6 In-Context Learning
Few-shot learning improves LLM performance in many tasks, such as question answering Brown et al. (2020). There are many strategies for selecting few-shot examples. Previous work suggested that similarity to the example in question is effective Paraschiv et al. (2023); Liu et al. (2022); Zebaze et al. (2024) was borne out in our early experiments. We use Linq-Embed-Mistral Kim et al. (2024) to create embeddings, and retrieve similar examples from our training data. Figure 2 shows results for few-shot up to six, for 4 non-reasoning models. For toxicity detection, examples help only for GPT4o, and requires at least three examples to surpass zero-shot. For toxic span prediction, one example is enough to increase the performance for Llama models, whereas Sonnet requires at least two examples. For target group prediction, few-shot learning helps all models, and the improvements are more dramatic, e.g., from 0.48 to 0.64 for Llama8b with one example. However, apart from prediction of target group, LLMs still do not surpass the performance of fine-tuned PLMs.
@@ 4.7 Error Analysis
The best results reported above (accuracy; F1) are from the RoBERTa classifier for toxicity (0.79; 0.79), the BERT classifier for fine-grained target (0.76; 0.72) and course-grained target (0.80; 0.79); SpanBERT for toxic spans (F1 of 0.70). In general, LLMs do better at toxicity detection than they do on the other two tasks, and reasoning/CoT helps very little. Here we look into differences across social groups for toxicity and target group prediction, whether span detection is easier when the whole sentence is the toxic span, and performance differences on machine-generated text.
Differences across social groups. For the error analysis, we broke down the accuracy results into accuracy per target social group for toxicity and target group detection. The Figure 3 accuracy heatmap for toxicity detection shows wide variation across target groups, ranging from as low as 41% up to 100%. Apart from the categories for "other gender" and "other ethnicity," where most models have very high performance, all models show non-uniform performance across groups. In some cases, pairs of models correlate in their performance (e.g. Spearman correlations are 0.89 for GPT4o and Sonnet; 0.87 for o4-mini and Llama70b; 0.83 for Sonnet and Llama70b), but in no case model performance correlates well with data support.
We also found large differences in target group accuracy across target social groups (see Appendix Figure 6). For example, Llama8b has a 42% difference in accuracy for "black" versus "white." The models struggle in categories such as "other country" or "other gender," possibly because these categories are necessarily more heterogenous.
Span detection for sentences versus subsentence spans. As mentioned in Section 3.1, toxic spans sometimes consist of the entire sentence (35% of the test set). Figure 4 plots F1 for full sentence toxic spans on the x-axis by subsentence spans on the y-axis. Most models, including SpanBERT, have much higher performance on full-sentence toxic spans. Distilled Llama models are far worse at detecting toxic subsentence spans.

### v2

@@ 3.3 Human Annotated Subset
Table 2 shows the proportion of toxic data for higher- or lower-level target group in the human-annotated subset, along with the breakdown by target group. Overall, 47.89% of the data was labeled toxic, but varied greatly by target group. For example, 75% of the examples for the social group “black” are toxic, compared with 33% for “native american.” Most target groups are ethnicity-based (18 %); the least frequent target group is for “disability.” (3.57 %). There is a variety within each higher-group as well: 4.50 % of the data targets LGBTQ+ but only 1.02 % targets Man.
@@
The total number of examples with annotations of toxic span is 5,506 (47.89% of 11,498, per Table 2). Some examples have multiple spans, so there are a total of 7,065 spans. Before computing descriptive statistics on span tokens, we preprocessed the data by applying stemming, and merging the string pairs “ni**a” and “ni**er”; “retarded” and “retard”; “stupid” and “stupidity”. Mean length of spans was 1.91 word tokens (median 1, max 11). In 33% of cases, the span constituted the entire sentence. Only 1,334 of the 7K spans are unique. The histogram in Figure 1 shows that the distribution of span tokens occurring more than once is highly skewed. It lists the 25 most common span tokens, with “ni**er,” “f**k,” “stupid,” “idiot,” and “b**ch” at the top. Some words are specific to particular groups such as “black” or “kike” but others such as “kill” or “stupid” apply in general.
4 Benchmarking Experiments
@@
Model Accuracy Precision Recall F1 PerspectiveAPI 0.50 0.48 0.77 0.59 LlamaGuard 0.67 0.77 0.43 0.55 ShieldGemma 0.66 0.70 0.50 0.58 RoBERTa 0.79 0.76 0.82 0.79 BERT 0.79 0.77 0.79 0.78 GPT4o 0.75 0.70 0.85 0.77 GPT4o-TextGrad 0.74 0.67 0.86 0.76 Sonnet 0.77 0.72 0.83 0.77 Sonnet-TextGrad 0.75 0.72 0.77 0.75 Llama70b 0.77 0.71 0.86 0.78 Llama70b-TextGrad 0.75 0.69 0.86 0.76 Llama8b 0.73 0.66 0.88 0.76 Llama8b-TextGrad 0.70 0.70 0.62 0.66 Reasoning and CoT D. Llama70b 0.75 0.69 0.84 0.76 D. Llama8b 0.70 0.63 0.90 0.74 o4-mini 0.78 0.71 0.90 0.79 GPT4o-cot 0.74 0.73 0.71 0.72 Llama8b-cot 0.73 0.67 0.84 0.75 Llama70b-cot 0.75 0.72 0.75 0.74 Sonnet-cot 0.73 0.73 0.67 0.70
Table 3: Toxicity detection results. D. is for distilled.
@@
Benchmarking is performed with the human-annotated data of 11,498 examples, randomly divided into validation (N=495), test (N=982) and training. Because the validation set is small, we constrained each target group to have a minimum of five examples. We compare performance of multiple baselines on prediction of toxicity, target social group, and for the toxic examples, prediction of spans. We report accuracy, precision, recall and F1, but omit accuracy for spans. We first compare PLMs, LlamaGuard, ShieldGemma and PerspectiveAPI with zero-shot LLMs on each task in turn. We also compare zero-shot LLMs with prompts from automated prompt-engineering. Finally, we test reasoning models and in-context learning.
@@
PLM baselines for toxicity and target social group use BERT devlin2019bert and RoBERTa liu2019robertarobustlyoptimizedbert with linear classifier layers, and for span prediction we use SpanBERT joshi2020spanbert (see appendix B). We include four LLMs: GPT4o openai2024gpt4ocard, Claude 3.7 Sonnet claude3.7, and Llama3.1 (70b and 8b) grattafiori2024llama3herdmodels (see Appendix E for the prompt). We use temperature=0, and set the seed for determinism, (but cf. atil2025nondeterminismdeterministicllmsettings). For toxicity, we include PerspectiveAPI lees2022newgenerationperspectiveapi, a neural network that provides a probability of toxicity, and ShieldGemma zeng2024shieldgemma and LlamaGuard inan2023llama, fine-tuned versions of Gemma-24b and Llama8b developed for content moderation. We used the validation set to identify the best probability for the binary class cutoff for PerspectiveAPI, which was 0.20.
@@ 4.1 Toxicity
The toxicity results in Table 3 show that fine-tuned PLMs perform slightly better than the LLMs. Except for Llama8b, LLMs perform similarly. Surprisingly, PerspectiveAPI’s accuracy is random. Both LlamaGuard and ShieldGemma underperform, indicating they do not generalize well. All models except LlamaGuard and ShieldGemma have higher recall which is preferable here, where false negatives are worse than false positives.
Model Accuracy Precision Recall F1 BERT 0.76 0.68 0.82 0.72 RoBERTa 0.73 0.63 0.81 0.70 GPT4o 0.75 0.67 0.78 0.70 GPT4o-TextGrad 0.71 0.57 0.69 0.61 Sonnet 0.74 0.62 0.75 0.67 Sonnet-TextGrad 0.75 0.63 0.68 0.62 Llama70b 0.65 0.55 0.68 0.58 Llama70b-TextGrad 0.69 0.34 0.38 0.34 Llama8b 0.48 0.15 0.15 0.15 Llama8b-TextGrad 0.43 0.10 0.11 0.10 Reasoning and CoT o4-mini 0.72 0.58 0.63 0.59 D. Llama70b 0.69 0.24 0.27 0.25 D. Llama8b 0.61 0.17 0.17 0.17 GPT4o-cot 0.75 0.68 0.75 0.70 Llama8b-cot 0.40 0.07 0.06 0.06 Llama70b-cot 0.68 0.50 0.58 0.53 Sonnet-cot 0.75 0.62 0.70 0.65
@@
Model Accuracy Precision Recall F1 BERT 0.80 0.77 0.82 0.79 RoBERTa 0.78 0.74 0.81 0.77 GPT4o 0.66 0.64 0.65 0.64 GPT4o-TextGrad 0.64 0.44 0.44 0.44 Sonnet 0.71 0.65 0.70 0.66 Sonnet-TextGrad 0.67 0.62 0.63 0.62 Llama70b 0.57 0.46 0.48 0.44 Llama70b-TextGrad 0.42 0.35 0.30 0.32 Llama8b 0.49 0.07 0.07 0.07 Llama8b-TextGrad 0.53 0.06 0.07 0.06 Reasoning and CoT o4-mini 0.67 0.45 0.47 0.45 D. Llama70b 0.64 0.17 0.18 0.17 D. Llama8b 0.56 0.07 0.07 0.07 GPT4o-cot 0.68 0.56 0.61 0.58 Llama8b-cot 0.44 0.05 0.05 0.05 Llama70b-cot 0.57 0.31 0.31 0.31 Sonnet-cot 0.70 0.55 0.60 0.57
@@ 4.2 Target Social Group
Following zampieri-etal-2023-target, we train two target group prediction baselines by fine-tuning BERT or RoBERTa with a linear neural classifier head. We also evaluate the same SOTA LLMs (prompts in Appendices F and G). Again, fine-tuned PLMs slightly outperform LLMs (cf. Table 4 for fine-grained results). Among LLMs, GPT4o and Sonnet outperform Llama models, with a particularly low F1 for Llama8b. Table 5 reports higher-level target results, where LLMs perform worse than on fine-grained targets (e.g., a 9% drop for GPT4o and 3% for Sonnet). Confusion matrices for GPT4o and Sonnet show confusions: GPT4o mixes “other” and “ethnicity” predictions with “no target,” Sonnet mixes “no target” with “ethnicity,” and both confuse “ethnicity” with “other.” This suggests that “other” as a high-level target is more ambiguous than among fine-grained labels, and that finer target granularity improves LLM performance.
(a) Toxicity.
(b) Fine-Grained Target Group.
(c) High-Level Target Group.
(d) Toxic Span.
Figure 2: Few-Shot Comparison for Zero to 6-Shot across Four Tasks with Four Models.
Model Precision Recall F1 SpanBERT 0.72 0.71 0.70 GPT4o 0.55 0.79 0.65 GPT4o-TextGrad 0.67 0.35 0.46 Sonnet 0.66 0.45 0.53 Sonnet-TextGrad 0.70 0.32 0.44 Llama70b 0.66 0.22 0.33 Llama70b-TextGrad 0.74 0.11 0.19 Llama8b 0.48 0.48 0.48 Llama8b-TextGrad 0.74 0.10 0.17 Reasoning and Cot o4-mini 0.63 0.40 0.49 D. Llama70b 0.45 0.87 0.59 D. Llama8b 0.43 0.58 0.40 Sonnet-cot 0.63 0.5 0.56 GPT4o-cot 0.55 0.63 0.59 Llama70b-cot 0.68 0.46 0.54 Llama8b-cot 0.68 0.16 0.26
@@
Notably, target-group information does not improve toxicity detection, as shown in Table 9. For RoBERTa/BERT models, we compared adding target information at the text level (“The target social group is <social group>”) versus at the embedding level. For LLMs, we tested two prompting variants: (1) assigning the target persona, motivated by evidence that persona plus self-correction can help xu-etal-2024-walking; and (2) explicitly including the target group in the prompt text.
@@ 4.3 Toxic Span
The span prediction results in Table 6 show that SpanBERT reaches 0.70 F1, outperforming the LLMs, and has balanced recall and precision. GPT4o’s F1 approaches SpanBERT’s, but the other LLMs do much worse.
4.4 Automated Prompt Engineering
Automated prompt optimization (APO) has been effective on code optimization, agent planning, mathematical reasoning, etc. yuksekgonul2025optimizing; ramnath2025systematicsurveyautomaticprompt. We used the SOTA TextGrad framework yuksekgonul2025optimizing, which uses an optimizer LLM to criticize the current prompt and suggest improvements; a new prompt is selected based on performance on the validation set. We used GPT4o as the optimizer for 15 iterations. Our training and validation data have 100 and 200 examples, respectively. Tables 3-6 show that TextGrad optimizations do not produce improvements. One potential reason for this is that we have already done manual prompt engineering, so they are close to optimal. Another reason might be that TextGrad is not very effective on subjective tasks.
@@ 4.5 Reasoning Models
Although reasoning helps in science or logic jaech2024openai; deepseekai2025deepseekr1incentivizingreasoningcapability; zhang2024chain; wei2023chainofthoughtpromptingelicitsreasoning, Tables 3-6 show no improvement on detection of toxicity or target group, and mixed results on span prediction. We experimented with chain-of-thought (CoT) wei2023chainofthoughtpromptingelicitsreasoning and the reasoning models o4-mini and R1 Distilled Llama70b/8b. CoT helps Llama70b and Sonnet, but not Llama8b and Gpt4o. Reasoning usually increases span prediction recall.
@@ 4.6 In-Context Learning
Few-shot learning improves LLM performance in many tasks, such as question answering brown2020language. Prior work—and our early experiments—suggested selecting demonstrations by similarity to the test example is effective paraschiv2023offensive; liu-etal-2022-makes; zebaze2024incontextexampleselectionsimilarity. We therefore use Linq-Embed-Mistral embeddings LinqAIResearch2024 to retrieve similar training examples. Figure 2 reports zero-to 6 shots for four non-reasoning models. For toxicity detection, examples help only GPT4o, and only with at least 3 examples. For toxic span prediction, one example improves Llama models, while Sonnet needs at least two. For target group prediction, few-shot learning benefits all models, with larger gains (e.g., Llama8b improves from 0.48 to 0.64 with one example). However, except for target group prediction, LLMs remain inferior to PLMs.
@@ 4.7 Error Analysis
To investigate factors behind our results, we look into four issues: 1) differences across targets for toxicity and target prediction; 2) whether span detection is better when the span is the whole sentence; 3) human vs. machine-generated text; 4) generalization of models trained on TRuST.
Differences across social groups. A breakdown of toxicity accuracy by target group revealed no patterns. Although most models performed well for the empty categories "other gender" and "other ethnicity," performance is otherwise highly variable across targets and models. A heatmap of toxicity detection for targets (cf. Figure 6 in Appendix) showed substantial variation across targets, e.g, 41%–100%. Some model pairs had correlated performance (Spearman $\rho=0.89$ for GPT4o–Sonnet; $0.87$ for o4-mini–Llama70b; $0.83$ for Sonnet–Llama70b), but no model’s performance correlated well with data support. Target group prediction also had large disparities across social groups (Figure 5 in Appendix L). Llama8b had a 42% accuracy gap between "black" and "white." Models also struggle in categories such as "other country" or "other gender," possibly because these groups are inherently more heterogeneous.
Span detection for sentences. As mentioned in Section 3.2, toxic spans sometimes consist of the entire sentence (35% of the test set). Most models, including SpanBERT, have much higher performance on full-sentence toxic spans. For example, SpanBERT has a score of 0.67 on subsentence spans, whereas has a score of 0.76 on full sentence. Distilled Llama models are far worse at detecting toxic subsentence spans. (Cf. Figure 3 in Appendix B, a plot of F1 for full sentence spans on the x-axis by subsentence spans on the y-axis.)
@@
Generalization of models trained on TRuST. To test generalization of models trained on TRuST, we trained PLM classifiers on the train subset of the manually annotated data (RoBERTa), or all the training data including synthetic labels (RoBERTa+), and tested on the OpenAI Moderation Dataset markov2023holistic. Table 7 shows that models trained on TRuST generalize well: their performance here is similar to performance on our test set. They outperform ShieldGemma, whereas LlamaGuard is the top performer.
Model Accuracy F1
RoBERTa 0.79 0.75
ShieldGemma 0.78 0.72
LlamaGuard 0.87 0.85
RoBERTa+ 0.82 0.79
Table 7: Toxicity detection results on OpenAI Moderation.
4.8 Automatic Labels in TRuST
The best results reported above (accuracy; F1) are from the RoBERTa classifier for toxicity (0.79; 0.79), the BERT classifier for fine-grained target (0.76; 0.72) and course-grained target (0.80; 0.79); SpanBERT for toxic spans (F1 of 0.70). These methods were used to annotate all remaining examples in TRuST, apart from the human annotation. The inter-annotator scores between the best PLMs and humans at 0.58/0.70/0.45 are close to human agreement, indicating automated label quality approaches human annotation quality.